OpenStudy (anonymous):

Can someone help me with this initial value problem? d^2y over dt^2 = 2/t^3. Given is: dr/dt = t=1 and that whole thing = 1. Another Given is: r(1) = 1. I have no idea how to go about this problem after finding the antiderivative for the d^2r over dx^2.

OpenStudy (freckles):

so you never meant to say anything about y right?

OpenStudy (anonymous):

Sorry meant r not y.

OpenStudy (freckles):

\[r''=2t^{-3}\]

OpenStudy (freckles):

So you said you integrated both sides: \[r'=2 \cdot \frac{t^{-3+1}}{-3+1}+C_1=2 \cdot \frac{t^{-2}}{-2}+C_1=-1 t^{-2}+C_1\] => \[r'=\frac{-1}{t^2}+C_1\]

OpenStudy (freckles):

So we are given r'(1)=1?

OpenStudy (anonymous):

Just r(1)=1

OpenStudy (freckles):

dr/dt = t=1? r'=t=1 I look at that is when t=1 ,r '=1

OpenStudy (freckles):

\[1=\frac{-1}{1^2}+C_1 \] Solve for C_1

OpenStudy (anonymous):

where does the r' come from?

OpenStudy (anonymous):

Oh! so another way to look at d^2r/dt^2 is r''?

OpenStudy (freckles):

yep yep

OpenStudy (anonymous):

Okay! Also just to clarify, when substituting t and r, just plug in one in newly acquired derivative where r and t are respectively? If r(1)=1 and t(1)=2, then plug in respectively and then ... what would happen?

OpenStudy (freckles):

well if you are given r'(a)=b then you use the function for r' If t=a then r'=b ------------------------ and if you are given r(c)=d then use the function r If t=c, then r=d You will use both of this conditions to find your constants

OpenStudy (anonymous):

So, since r(1)=1, then t(1)=2, then r of t = 1? Somehow I confused myself...

OpenStudy (freckles):

t(1)=2 makes no sense

OpenStudy (freckles):

I'm sorry to be so blunt

OpenStudy (anonymous):

No, it makes no sense to me either.

OpenStudy (freckles):

You are given r'(1)=1 and r(1)=1

OpenStudy (freckles):